OpenStudy (anonymous):

Please help :0 Find the term containing x^42 of (x^3-3)^16

OpenStudy (anonymous):

well first we need \(\frac{42}{3}=14\) to know which term we want

OpenStudy (anonymous):

so it is either the 14th term or the second term depending on how you are counting. makes no difference. the term looks like \[\dbinom{16}{2}(x^3)^{14}(-3)^2\]

OpenStudy (anonymous):

\[\dbinom{16}{2}=\frac{16\times 15}{2}=8\times 15=120\] and \((-3)^2=9\) multiply to get your answer

OpenStudy (anonymous):

wow my drawing really sucked.... oh well!!

OpenStudy (anonymous):

Thanks so much satellite73!